OpenStudy (anonymous):

Find the product 12times1/3times(-1)times(-1/2) A.18 B. 2 C.-2 D.-2 1/6

OpenStudy (anonymous):

@pierodog

OpenStudy (anonymous):

or anyone

OpenStudy (anonymous):

@Jackky

OpenStudy (anonymous):

@pierodog @Jackky

OpenStudy (misty1212):

HI!

OpenStudy (anonymous):

hi I need help

OpenStudy (misty1212):

is it this \[12\times \frac{1}{3}\times (-1)\times (-\frac{1}{2})\]??

OpenStudy (anonymous):

yes @misty1212

OpenStudy (misty1212):

ok lets take it nice and slow how many minus signs you see?

OpenStudy (anonymous):

2

OpenStudy (misty1212):

right and you know that makes it positive right?

OpenStudy (anonymous):

yeah

OpenStudy (misty1212):

ok now we have a choice, we can do two at a time or just multiply straight across

OpenStudy (anonymous):

multiply straight across

OpenStudy (misty1212):

straight across it looks like \[\frac{12\times 1\times 1}{2\times 3}\]

OpenStudy (anonymous):

yes that one

OpenStudy (misty1212):

ok then we multiply and get \[\frac{12}{6}\] and what is 12 divided by 6?

OpenStudy (anonymous):

2

OpenStudy (jackky):

wait wouldnt it be A though

OpenStudy (misty1212):

that is the final answer

OpenStudy (jackky):

OH ok thank you i had that thought

OpenStudy (anonymous):

tnx so much but I need help in grammer

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]